Where 

C=250+0.75Yd

I=350

G=800

X=500

M=200+0.4Y

T=0.15Y


Use the above information to answer the following questions 


1) if C=250+0.75Yd interpret the meaning of the value 250

2)state the values of MPT,MPC and MPS on disposable income and MPM

3)find MPC and MPS on national income 

4)find the level of national income in the economy 

5) calculate net exports and interpret 

6) if investment rises by 15% ,what will be the change in National income 


1
Expert's answer
2022-01-25T06:46:49-0500

Solution:

1.). 250 is autonomous consumption.

Autonomous consumption is defined as consumer expenditures that must be made even when there is no disposable income. That is, it happens when income levels are zero.

 

2.). MPC = 0.75

MPS = 1 – 0.75 = 0.25

MPT = "\frac{-0.25}{(1 - 0.25)} = -0.33"

MPM = 0.4

 

3.). MPC = 0.75

MPS = 1 – 0.75 = 0.25

 

4.). National Income (Y): AD = AS

Y = C + I + G + X – M

C = 250 + 0.75Yd = 250 + 0.75 (Y – T) = 250 + 0.75 (Y – 0.15Y) = 250 + 0.6375Y

Y = 250 + 0.6375Y + 350 + 800 + 500 – 200 + 0.4Y

Y – 0.6375Y + 0.4Y = 1700

0.7625Y = 1700

Y = 2229.51

National Income (Y) = 2,229.51

 

5.). Net exports = Exports (X) – Imports(M)

500 – 200 + 0.4Y = 500 – 200 + 0.4(2,229.51) = 500 – 200 + 891.81 = 500 – 1091.81 = (591.81)

Net Exports = (591.81)

 

6.). Investment = 350

New investment = 350 + 15% = 402.5

Y = 250 + 0.6375Y + 402.5 + 800 + 500 – 200 + 0.4Y

Y – 0.6375Y + 0.4Y = 1752.5

0.7625Y = 1752.5

Y = 2,298.36

New national Income (Y) = 2,298.36

The change in national income = 2,298.36 – 2,229.51 = 68.85